OpenStudy (anonymous):

You collected the following data about the color of dresses at the prom: Red Black Blue white other 14 37 22 56 33 When making a pie chart for this data, what percentage of the circle would the sector labeled "Red" take up? Round your answer to the nearest whole percent. A. 9% B. 14% C. 23% D. 25%

OpenStudy (anonymous):

14% right?

OpenStudy (austinl):

\(\displaystyle \frac{\text{# red dresses}}{\text{total # of dresses}}\)

OpenStudy (anonymous):

14/162 do i divide then?

OpenStudy (anonymous):

14% would be correct if there were 100 colored dresses only, but there are more than 100 dresses.

OpenStudy (austinl):

Yep, \(\displaystyle \frac{14}{14+37+22+56+33}=~?\%\)

OpenStudy (anonymous):

it would still be 14% correct?

OpenStudy (austinl):

Nope :( \(\displaystyle \frac{14}{14+37+22+56+33}=\frac{14}{162}\times 100=~?\%\)

OpenStudy (anonymous):

=8.641975....

OpenStudy (anonymous):

so 9%

OpenStudy (austinl):

It is indeed :)

OpenStudy (anonymous):

good job!

OpenStudy (austinl):

You as well lol!